The velocity v in a fluid in a two dimensional model at any point (x,y) in the flow is given by the function below.


v=4x2y−32xy+48y−24x

Find the stationary points for this function

(I)

Find the x coordinate of the stationary point with the larger x coordinate. Give your answer to AT LEAST FOUR PLACES OF DECIMALS. Type in the number only ie leave out the "x=" (4 marks)



This question accepts numbers or formulas.
Plot | Help | Switch to Equation Editor | Preview

(II)

Find the corresponding y coordinate of the stationary point whose x coordinate you have given in (i). Give your answer to AT LEAST FOUR PLACES OF DECIMALS. Type in the number only ie leave out the "y=" (4 marks)



This question accepts numbers or formulas.
Plot | Help | Switch to Equation Editor | Preview

(III)

What is the nature of this stationary point?
Saddle (enter 1)
Maximum (enter 2)
Minimum (enter 3) (2 marks)



This question accepts numbers or formulas.
Plot | Help | Switch to Equation Editor | Preview

(IV)

Find the value of the velocity at this stationary point. Give your answer to AT LEAST FOUR PLACES OF DECIMALS. Type in the number only ie leave out the "v=" (2 marks)



This question accepts numbers or formulas.
Plot | Help | Switch to Equation Editor | Preview

(V)

Find the x coordinate of the stationary point with the smaller x coordinate. Give your answer to AT LEAST FOUR PLACES OF DECIMALS. Type in the number only ie leave out the "x=" (4 marks)



This question accepts numbers or formulas.
Plot | Help | Switch to Equation Editor | Preview

(VI)

Find the corresponding y coordinate of the stationary point whose x coordinate you have given in (v). Give your answer to AT LEAST FOUR PLACES OF DECIMALS. Type in the number only ie leave out the "y=" (4 marks)



This question accepts numbers or formulas.
Plot | Help | Switch to Equation Editor | Preview

(VII)

What is the nature of this stationary point?
Saddle (enter 1)
Maximum (enter 2)
Minimum (enter 3) (2 marks)



This question accepts numbers or formulas.
Plot | Help | Switch to Equation Editor | Preview

(VIII)

Find the value of the velocity at this stationary point. Give your answer to AT LEAST FOUR PLACES OF DECIMALS. Type in the number only ie leave out the "v=" (2 marks)
Answer #1

V(x,y) = 4*?y - 32xy +494 - 24% Ver de = fxy – 32y – 24 Bu y = 4x² – 32x tag Vy - To get stationary points Vx= 0 ? Vy- o Sxy-(6,3/2) NXX = $(3/2) = 12 ; Vyy=0; Vey = 8(6) -32 = 16 H = (12)(0) - (163 = -25540. - H2O =) (6, 7) is a seedle point (1) x =